OpenStudy (anonymous):

Using the sum formula, show cos(x+pi)=-cos(x) Can someone tell me the steps of what to do. I can do the sum formula but isn't it different since it equals -cos(x)?

OpenStudy (anonymous):

\[\cos(\alpha +\beta)=\cos(\alpha)\cos(\beta)-\sin(\alpha)\sin(\beta)\] put \(\alpha =x, \beta=\pi\) and your answer will drop right out

OpenStudy (anonymous):

So just use the formula. Theres no difference even though it equals -cos(x)? Like i don't do the even/odd test after?

OpenStudy (anonymous):

no make a direct substitution the formula is the formula

OpenStudy (anonymous):

the question says "SHOW \(\cos(x+\pi)=-\cos(x)\)" make that sub and you will get that answer

OpenStudy (anonymous):

got it. So cos(-x)?

OpenStudy (anonymous):

or -cos(x)